OpenStudy (anonymous):

HELP ME PLEASE!!! A total of 440 tickets were sold for the school play. They were either adult tickets or student tickets. There were 60 fewer student tickets sold than adult tickets. How many adult tickets were sold?

OpenStudy (anonymous):

I belive but not sure that it could be 280

OpenStudy (anonymous):

that's not the answer but its close.

OpenStudy (anonymous):

what i did is divide it by two and then on one side minus 60 and add it to the other side

OpenStudy (anonymous):

a + s = 440 s = a - 60 now sub a - 60 in for s and solve for a

OpenStudy (anonymous):

can someone help me with another question? Gabrielle's age is three times Mikhail's age. The sum of their ages is 104 . What is Mikhail's age?

OpenStudy (anonymous):

g = 3m g + m = 104 now sub 3m in for g in the 2nd equation and solve for m

OpenStudy (anonymous):

okay, I got 26

OpenStudy (anonymous):

Get adult tickets = a student tickets = s And find an equation for them : a+s=440 And you said : There were 60 fewer student tickets sold than adult tickets. So can say : s=a-60 Then get a-60 to s . So you have : a-60+a=440 2a=440-60 Find a .

OpenStudy (anonymous):

26 is correct

OpenStudy (anonymous):

Mikhail's age=m Gabrielle's age = g We have : g=3m And : m+g = 104 m+3m=104 4m=104 m=26

OpenStudy (anonymous):

thank you e.ali! as well.

OpenStudy (anonymous):

In this question just get a thing like words for each things and write an equation for them . And remember your equation is very important .

OpenStudy (anonymous):

And welcome.